OpenStudy (anonymous):

3x/2 - x = x/6 - 4/3 Solve

OpenStudy (anonymous):

\[\frac{ 3x }{ 2 } - x - \frac{ x }{ 6 } = \frac{ -4 }{ 3 }\]

OpenStudy (anonymous):

How do you solve it

OpenStudy (anonymous):

Take lcm of 2,6, 1 (i.e 12) \[\frac{ 6*(3x) }{ 12 } - \frac{ 12*x }{ 12 } - \frac{ 2*x }{ 12 } = \frac{ -4 }{ 3 }\] \[\frac{ 18x - 12x - 2x }{ 12 } = \frac{ -4 }{ 3 }\] solve for x

OpenStudy (anonymous):

The answer is -4/3

OpenStudy (anonymous):

Is this right

OpenStudy (anonymous):

\[\frac{ 4x }{ 12 }= \frac{ -4 }{ 3 }\] \[ 4x = \frac{ -4 *12}{ 3 }\] \[ x = \frac{ -4 *12}{ 3 *4 }\] \[ x = \frac{ -1 *12}{ 3 }\] \[ x = \frac{ -1 *4}{ 1 }\] \[ x = -4\]

OpenStudy (anonymous):

Alright I see how to do it